OpenStudy (anonymous):

Evaluate the following limit (in order to find a tangent line adjacent to the function). See problem below...

OpenStudy (anonymous):

\[\ \Huge Evaluate: lim_{h \rightarrow 0} \frac{\sqrt{1+h} - \sqrt{1}}{h}.\]

OpenStudy (dumbcow):

multiply by conjugate on top and bottom \[\frac{\sqrt{1+h}-1}{h}*\frac{\sqrt{1+h}+1}{\sqrt{1+h}+1}\]

OpenStudy (anonymous):

Okay so when I do that, @dumbcow, I end up with 1 is that correct?

OpenStudy (dumbcow):

no...how did you get 1? FOIL the top

OpenStudy (anonymous):

I did

OpenStudy (dumbcow):

\[(\sqrt{h+1})^{2} +\sqrt{h+1}-\sqrt{h+1} -1 = h\]

OpenStudy (anonymous):

1/2 does seem right either... I got h on the numerator

OpenStudy (dumbcow):

yes the limit is 1/2